The total market value of the equity of Okefenokee Condos is $8 million, and the total value of its debt is $2 million. The treasurer estimates that the beta of the stock currently is 0.6 and that the expected risk premium on the market is 10%. The Treasury bill rate is 5%, and investors believe that Okefenokee's debt is essentially free of default risk.

a. What is the required rate of return on Okefenokee stock? (Do not round intermediate calculations. Enter your answer as a whole percent.)

b. Estimate the WACC assuming a tax rate of 21%. (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)

c. Estimate the discount rate for an expansion of the company’s present business. (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)

d. Suppose the company wants to diversify into the manufacture of rose-colored glasses. The beta of optical manufacturers with no debt outstanding is .8. What is the required rate of return on Okefenokee’s new venture? (You should assume that the risky project will not enable the firm to issue any additional debt.) (Do not round intermediate calculations. Enter your answer as a whole percent.)
Answer #1

Given,

Equity = $ 8 million

Debt = $ 2 million

Company value = $ 8 million + $ 2 million = $ 10 million

Risk free rate = 5% or 0.05

Beta of the stock = 0.6

Risk premium on the market = 10% or 0.10

Solution :-

(a) Requiredl retuan tPlm-A whtre Bisk frce rate ot the Stock Beta Risk paemi Cn the market 11 Nows Required retuan =0.05 + 0Sy(1-0.31 3.95 2 S% (0.79) 1 Cost of equity (ke) No equity XKe (ompary Vatue De bt Company Value WACC $8 million $10 millionBeta 0.8 N Cost of equity ke)Xe t B nisk Paemium) O.os + 0.8(0.1o) = 0.05+ 0.08 =0.13 13/% eguired aehun an Okefenokcels ncw
Answer #2

d) Required rate of return= treasury bill rate + new beta (expected risk premium)

=5%+0.8(10%)

=13%

answered by: Chloe Cervera